OpenStudy (anonymous):

Which of the following represents seventh root of x to the sixth power in exponential form?

OpenStudy (anonymous):

\[\large \sqrt[n]{x^m}=x^{\frac{m}{n}}\]

OpenStudy (anonymous):

thank you!!!!